OpenStudy (anonymous):

CHECK MY ANSWER PLEASE? Segment RT has the endpoints (–3, 4) and (11, 2). What is its midpoint? (1, 1) (4, 3) my answer ♥ (2, 2) (1, 3)

OpenStudy (anonymous):

-2+6 5 + (-9) 4 -4 ----- , --------- = ---- , ------ 2 2 2 2 (2, -2)

OpenStudy (anonymous):

x1 + x2 y1 + y2 mid point = -------- , -------- 2 2

OpenStudy (anonymous):

the answer is 2,2

OpenStudy (ranga):

To find mid point, add the x coordinates and divide by 2. Add the y coordinates and divide by 2. (-3 + 11)/2 = 8/2 = 4 (4 + 2)/2 = 6/2 = 3 So (4,3) is the correct answer and you have it right.